# Spin-Orbit Misalignments of Eccentric Black Hole Mergers in AGN Disks

### v1

@@ 4.4 Retrograde orbiters and warp radius
The alignment of isolated binary-disk systems (as opposed to disk-embedded binaries treated here) is typically governed by an analog to the Bardeen-Petterson effect whereby differentially precessing annuli, whose angular momentum vectors are tilted relative to that of the binary, are viscously damped into the plane of the binary orbit (Bardeen & Petterson, 1975; Nixon et al., 2011). [^3] Provide the disk is sufficiently viscous. This is effective out to the warp radius,
@@
where the viscous time characterized by vertical sheer viscosity $\nu_{2}$ becomes equal to the local precession time (Pringle, 1992; Scheuer & Feiler, 1996; King et al., 2005). For disk-embedded binaries, if the accretion flow were sufficiently independent from the large scale shear of the AGN-disk, the associated timescale could alternatively be applied to estimate $\tau_{\rm align}$.
@@ 5 Caveats
  • Super-Eddington accretion: Our prescription assumes accretion rates that are often highly super-Eddington, raising the question of whether such flows can be sustained in AGN disks. If that is not the case, alignment efficiencies would be significantly reduced, as $\dot{M}$ would occur at a much slower rate, unless $f_{\rm rot}$ was raised commensurately. There remains substantial uncertainty as feedback from radiation pressure and jet-driven outflows could strongly regulate or even suppress such accretion (Pan & Yang, 2021; Tagawa et al., 2022). It is therefore important to understand whether AGN disks can sustain these regimes, since the physical justification of large alignment efficiencies (e.g. $f_{\rm rot}\gtrsim 1$) would also depend on the availability of super-Eddington accretion. If the actual accretion rate is much lower than assumed in our models, this would imply that many of our circular binaries would also not have time to align, and skew the anticipated $\chi_{\rm eff}$ distribution of the AGN channel more towards zero. Whether sustained super-Eddington accretion is possible in AGN disks is still an unresolved problem, and thus our treatment still represents one plausible scenario.
  • Scattering in and out of the disk: In our alignment prescription we assume that binaries remain confined to the disk after formation. In reality, repeated encounters can scatter black holes out of the disk plane, temporarily removing them from the dense gaseous environment where alignment operates most efficiently. Moreover, as argued by Ford & McKernan (2025), the deep SMBH potential well typically retains BHs even after they are scattered out of the disk, causing them to remain bound and re-cross the disk on orbital timescales. The net impact on the inclination distribution therefore depends on how often and on which timescales systems leave and re-cross the disk.
@@
  • Turbulence: throughout the paper we neglect the effect of turbulence. Turbulence in AGN disks can yield stochasticity in the accretion process and randomize angular momentum transport (Chen & Lin, 2023; Mishra & Calcino, 2024), effectively reducing the efficacy of alignment (Trani & Di Cintio, 2025). In this case, even binaries embedded in the densest regions of the disk might experience weaker alignment than predicted. The details on strength effect of turbulence in AGN disks and its role in shaping the evolution of BBHs remain an open problem in astrophysics.
@@
  • Repeated interactions and hierarchical mergers: In our $N$-body simulations, the integration is terminated once the binary enters the inspiral phase. The majority of these mergers are three-body mergers, in which the third companion remains bound to the system at the time of merger assembly. This raises the possibility of further dynamical interactions: while the binary is inspiraling and potentially re-aligning through accretion, the bound single may return and perturb the system again. Even if the single is temporarily ejected beyond the Hill sphere during the initial encounter, the presence of the AGN disk favours its re-formation (e.g. Rowan et al., 2025b), enabling subsequent close passages. Most importantly, the single is not only capable of re-entering the system and scattering the binary during its inspiral, but it may also merge with it, leading to a hierarchical merger (Samsing & Ilan, 2018, 2019). Recent studies have even suggested that distinct gravitational wave events could be connected through their proximity in sky localization, raising the possibility that one merger directly followed another within the same AGN environment (Veske et al., 2020; Li & Fan, 2025). Such repeated scatterings during the inspiral phase could alter the orbital configuration of the merging binary, affecting the realignment process. Although our present simulations do not follow this process beyond the initial inspiral, exploring the combined role of repeated encounters, gas-assisted recapture, and the possibility of hierarchical mergers will examined future work.
@@ 6 Conclusions
  • Correlation between eccentricity and effective spin in the AGN channel: Although $\theta_{B}$ is not directly measurable, if gas aligns the individual BH spins with the disk angular momentum (e.g. Tagawa et al., 2020b; Vajpeyi et al., 2022), then $\chi_{\rm eff}$ follows $\cos\theta_{B}$ (e.g. Samsing et al., 2022), allowing us to link orbital tilt to the effective spin. We find that the skewness of $\chi_{\rm eff}$ depends strongly on eccentricity and alignment efficiency. Quasi-circular binaries ($e_{\rm 10Hz}\lesssim 0.01$) consistently re-align and yield positive $\chi_{\rm eff}$. Eccentric binaries, on the other hand, can retain isotropy unless alignment is sufficiently effective: at our fiducial value $f_{\rm rot}=0.1$, the distribution remains nearly flat, but for higher efficiencies ($f_{\rm rot}\gtrsim 1$) or denser disks, the skewness recovers a positive bias. This provides a natural way to map a correlation between binary eccentricity and $\chi_{\rm eff}$, where circular systems preferentially yield positive values while eccentric ones span a wider and potentially isotropic range. Furthermore, binaries that merge before achieving full alignment are expected to show enhanced in-plane spin components ($\chi_{P}$) (Ford & McKernan, 2025). In such cases, even when the total spin magnitude is large, the effective spin can remain modest, offering a complementary signature of the AGN channel. This framework naturally encompasses massive events such as GW190521, with its observed spin–orbit tilt and possible eccentricity (Abbott et al., 2020; Romero-Shaw et al., 2020, 2022; Gayathri et al., 2022), and GW231123, whose highly spinning components may have grown through accretion (The LIGO Scientific Collaboration et al., 2025). Both could plausibly have originated in AGN disks (Delfavero et al., 2025a; Bartos & Haiman, 2025). More broadly, AGN-assisted mergers may produce $\chi_{\rm eff}$ distributions ranging from strongly positive to nearly isotropic depending on merger properties, offering a direct observational test for future LVK analyses. This is consistent with current LVK results, which show $\chi_{\rm eff}$ values clustered around zero with a mild positive skew and only a small tail toward large spins (e.g. Abbott et al., 2023). A mixed population of mergers (circular, aligned binaries yield $\chi_{\rm eff}>0$ and eccentric, partially aligned systems remain near zero) in AGN disks can potentially explain this observed distribution.
@@ 6.1 Future Prospects
The central result of this work is the emerging correlation between binary eccentricity and the effective spin parameter $\chi_{\rm eff}$ in AGN disks. Our analysis shows that quasi-circular binaries are likely to realign efficiently with the disk and therefore yield positive or near-zero $\chi_{\rm eff}$, while eccentric binaries can retain substantial spin–orbit misalignments unless alignment is extremely efficient. This establishes eccentricity as a potential predictor of the $\chi_{\rm eff}$ distribution in the AGN channel and provides a direct observational test for current and future gravitational wave detections. Our framework represents a preliminary physical model for alignment in AGN disks. While not yet including the full complexity of binary–disk coupling, it captures the essential processes that determine whether binaries merge aligned or misaligned with the disk. A number of factors act to influence the effective alignment efficiency. Eccentric binaries shrink rapidly through GW emission, leaving less time for gas torques to act. Gas itself can accelerate orbital decay, as hydrodynamical simulations of binary–single encounters show that the compactness of triples decreases by up to two orders of magnitude in dense environments (e.g. Rowan et al., 2025b; Wang et al., 2025b, a). This further limits alignment, especially in regions of the disk with high density and strong accretion. Moreover, our present study has focused on BH–BH encounters, but the dominant interactions in AGN disks are likely between stars and stellar-mass BHs. Such star–BH scatterings can repeatedly perturb binaries during their inspiral, effectively altering the alignment process.
@@ Acknowledgments
GF gratefully acknowledges valuable discussions with the GW-Astro group at NBIA, in particular with Philip Kirkeberg and David O’Neill, the participants of the DYNAMIX meeting in Cambridge, as well as Alex Dittmann, Barry McKernan and Saavik Ford, whose comments provided important insights for this work. The Center of Gravity is a Center of Excellence funded by the Danish National Research Foundation under grant No. 184. The research leading to this work received funding from the Villum Fonden Grant No. 29466, ERC Starting Grant no. 101043143, the Independent Research Fund Denmark via grant ID 10.46540/3103-00205B, and by the European Union’s Horizon 2023 research and innovation program under Marie Sklodowska-Curie grant agreement No. 101148364. The Tycho supercomputer hosted at the SCIENCE HPC center at the University of Copenhagen was used for performing the presented simulations.

### v2

@@ 4.4 Retrograde orbiters and warp radius
The alignment of isolated binary-disk systems (as opposed to disk-embedded binaries treated here) is typically governed by an analog to the Bardeen-Petterson effect whereby differentially precessing annuli, whose angular momentum vectors are tilted relative to that of the binary, are viscously damped into the plane of the binary orbit (Bardeen and Petterson, 1975; Nixon et al., 2011). [^4] Provided the disk is sufficiently viscous. This is effective out to the warp radius,
@@
where the vertical shear viscosity $\nu_{2}$ is set equal to the standard $\alpha$-parametrized azimuthal viscosity, $\nu_{2}=\alpha\,c_{s}\,H$, computed from the local disk properties. This is a simple model for the structure of the accretion flow that does not account for local deviations from the ambient disk (e.g., Li et al., 2022b). At the warp radius, the viscous time associated to $\nu_{2}$ becomes equal to the local precession time (Pringle, 1992; Scheuer and Feiler, 1996; King et al., 2005). For disk-embedded binaries, if the accretion flow were sufficiently independent from the large scale shear of the AGN-disk, the associated timescale could alternatively be applied to estimate $\tau_{\rm align}$.
@@ 5 Caveats
  • Accretion onto the individual BBH components: A key assumption of our analysis is that only a limited fraction of the ambient disk gas flows into the binary, such that gas–binary coupling primarily manifests through gravitational torques rather than substantial BBH mass growth. Recent hydrodynamical simulations indicate that accretion onto embedded binaries is highly time-variable, with orbital evolution determined by complex gas dynamics and gravitational torques, rather than by a simple, steady accretion flow (e.g., Li and Lai, 2022, 2023). These studies further indicate that the gas distribution near the binary is strongly non-axisymmetric (Rowan et al., 2023) and therefore only a fraction of the surrounding gas might become bound to the individual black holes. If, however, circum-binary flows were found to be highly efficient at feeding the binary, the resulting growth of the black hole masses would shorten the gravitational-wave merger timescale (Eqs. (16)–(17)). In this regime, binaries could merge before alignment is completed, even for circular systems. While this does not affect our conclusion that eccentric binaries typically merge before alignment due to their short gravitational-wave timescales, it could reduce the fraction of aligned circular mergers and weaken the predicted correlation between eccentricity and $\chi_{\rm eff}$ in selected parts of the disk.
  • Super-Eddington accretion: Our prescription assumes accretion rates that are often highly super-Eddington, raising the question of whether such flows can be sustained in AGN disks. If that is not the case, alignment efficiencies would be significantly reduced, as $\dot{M}$ would occur at a much slower rate, unless $f_{\rm rot}$ was raised commensurately. There remains substantial uncertainty as feedback from radiation pressure and jet-driven outflows could strongly regulate or even suppress such accretion (Pan and Yang, 2021; Tagawa et al., 2022). It is therefore important to understand whether AGN disks can sustain these regimes, since the physical justification of large alignment efficiencies (e.g. $f_{\rm rot}\gtrsim 1$) would also depend on the availability of super-Eddington accretion. If the actual accretion rate is much lower than assumed in our models, this would imply that many of our circular binaries would also not have time to align, and shift the anticipated $\chi_{\rm eff}$ distribution of the AGN channel more towards zero. Whether sustained super-Eddington accretion is possible in AGN disks is still an unresolved problem, and thus our treatment still represents one plausible scenario.
  • Scattering in and out of the disk: In our alignment prescription we assume that binaries remain confined to the disk after formation. In reality, repeated encounters can scatter black holes out of the disk plane, temporarily removing them from the dense gaseous environment where alignment operates most efficiently. Moreover, as argued by Ford and McKernan (2025), the deep SMBH potential well typically retains BHs even after they are scattered out of the disk, causing them to remain bound and re-cross the disk on orbital timescales. The net impact on the inclination distribution therefore depends on how often and on which timescales systems leave and re-cross the disk.
@@
  • Turbulence: throughout the paper we neglect the effect of turbulence. Turbulence in AGN disks can yield stochasticity in the accretion process and randomize angular momentum transport (Chen and Lin, 2023; Mishra and Calcino, 2024), effectively reducing the efficacy of alignment (Trani and Di Cintio, 2025). In this case, even binaries embedded in the densest regions of the disk might experience weaker alignment than predicted. Turbulent accretion and supersonic motion relative to the disk can also misalign or flip mini-disks, potentially decoupling the individual black-hole spins from $\mathbf{J}_{\rm AGN}$ (Li et al., 2022a; Chen and Lin, 2023). The details on strength effect of turbulence in AGN disks and its role in shaping the evolution of BBHs remain an open problem in astrophysics.
@@
  • Repeated interactions and hierarchical mergers: In our $N$-body simulations, the integration is terminated once the binary enters the inspiral phase. The majority of these mergers are three-body mergers, in which the third companion remains bound to the system at the time of merger assembly. This raises the possibility of further dynamical interactions: while the binary is inspiraling and potentially re-aligning through accretion, the bound single may return and perturb the system again. Even if the single is temporarily ejected beyond the Hill sphere during the initial encounter, the presence of the AGN disk favours its re-formation (e.g. Rowan et al., 2025b), enabling subsequent close passages. Most importantly, the single is not only capable of re-entering the system and scattering the binary during its inspiral, but it may also merge with it, leading to a hierarchical merger (Samsing and Ilan, 2018, 2019). Recent studies have even suggested that distinct gravitational wave events could be connected through their proximity in sky localization, raising the possibility that one merger directly followed another within the same AGN environment (Veske et al., 2020; Li and Fan, 2025). Such repeated scatterings during the inspiral phase could alter the orbital configuration of the merging binary, affecting the realignment process. Although our present simulations do not follow this process beyond the initial inspiral, exploring the combined role of repeated encounters, gas-assisted recapture, and the possibility of hierarchical mergers will examined future work.
@@ 6 Conclusions
  • Correlation between eccentricity and effective spin in the AGN channel: Although $\theta_{B}$ is not directly measurable, if gas aligns the individual BH spins with the disk angular momentum (e.g. Tagawa et al., 2020a; Vajpeyi et al., 2022), then $\chi_{\rm eff}$ follows $\cos\theta_{B}$ (e.g. Samsing et al., 2022), allowing us to link orbital tilt to the effective spin. We find that the distribution of $\chi_{\rm eff}$ values depends strongly on eccentricity and alignment efficiency. Quasi-circular binaries ($e_{\rm 10Hz}\lesssim 0.01$) consistently re-align and yield positive $\chi_{\rm eff}$. Eccentric binaries, on the other hand, can retain isotropy unless alignment is sufficiently effective: at our fiducial value $f_{\rm rot}=0.1$, the distribution remains nearly flat, but for higher efficiencies ($f_{\rm rot}\gtrsim 1$) or denser disks, it recovers a positive bias. This provides a natural way to map a correlation between binary eccentricity and $\chi_{\rm eff}$, where circular systems preferentially yield positive values while eccentric ones span a wider and potentially isotropic range. Furthermore, binaries that merge before achieving full alignment are expected to show enhanced in-plane spin components ($\chi_{P}$) (Ford and McKernan, 2025). In such cases, even when the total spin magnitude is large, the effective spin can remain modest, offering a complementary signature of the AGN channel. This framework naturally encompasses massive events such as GW190521, with its observed spin–orbit tilt and possible eccentricity (Abbott et al., 2020; Romero-Shaw et al., 2020, 2022; Gayathri et al., 2022), and GW231123, whose highly spinning components may have grown through accretion (The LIGO Scientific Collaboration et al., 2025). Both could plausibly have originated in AGN disks (Delfavero et al., 2025a; Bartos and Haiman, 2025). More broadly, AGN-assisted mergers may produce $\chi_{\rm eff}$ distributions ranging from strongly positive to nearly isotropic depending on merger properties, offering a direct observational test for future LVK analyses. This is consistent with current LVK results, which show $\chi_{\rm eff}$ values clustered around zero with a mild positive average-weighted distribution and only a small tail toward large spins (e.g. Abbott et al., 2023). A mixed population of mergers (circular, aligned binaries yield $\chi_{\rm eff}>0$ and eccentric, partially aligned systems remain near zero) in AGN disks can potentially explain this observed distribution.
@@ 6.1 Future Prospects
The central result of this work is the emerging correlation between binary eccentricity and the effective spin parameter $\chi_{\rm eff}$ in AGN disks. Our analysis shows that quasi-circular binaries are likely to realign efficiently with the disk and therefore yield positive or near-zero $\chi_{\rm eff}$, while eccentric binaries can retain substantial spin–orbit misalignments unless alignment is extremely efficient. This establishes eccentricity as a potential predictor of the $\chi_{\rm eff}$ distribution in the AGN channel and provides a direct observational test for current and future gravitational wave detections. Our framework represents a preliminary physical model for alignment in AGN disks. While not yet including the full complexity of binary–disk coupling, it captures the essential processes that determine whether binaries merge aligned or misaligned with the disk. A number of factors act to influence the effective alignment efficiency. Eccentric binaries shrink rapidly through GW emission, leaving less time for gas torques to act. Gas itself can accelerate orbital decay, as hydrodynamical simulations of binary–single encounters show that the compactness of triples decreases by up to two orders of magnitude in dense environments (e.g. Rowan et al., 2025b; Wang et al., 2025a, b). This further limits alignment, especially in regions of the disk with high density and strong accretion. Moreover, our present study has focused on BH–BH encounters, but the dominant interactions in AGN disks are likely between stars and stellar-mass BHs. Such star–BH scatterings can repeatedly perturb binaries during their inspiral, effectively altering the alignment process.
@@ Acknowledgments
We gratefully acknowledge valuable discussions with Philip Kirkeberg and David O’Neill, the participants of the DYNAMIX meeting in Cambridge, as well as Alex Dittmann, Barry McKernan and Saavik Ford, whose comments provided important insights for this work. The authors also thank the anonymous referee for the insightful suggestions and comments. The Center of Gravity is a Center of Excellence funded by the Danish National Research Foundation under grant No. 184. The research leading to this work received funding from the Villum Fonden Grant No. 29466, ERC Starting Grant no. 101043143, the Independent Research Fund Denmark via grant ID 10.46540/3103-00205B, and by the European Union’s Horizon 2023 research and innovation program under Marie Sklodowska-Curie grant agreement No. 101148364. The Tycho supercomputer hosted at the SCIENCE HPC center at the University of Copenhagen was used for performing the presented simulations.